OpenStudy (anonymous):

What is the slope of the line that passes through the points (−5,4) and (15,−4)?

OpenStudy (nurali):

(-5,4),(15,-4) x1 = -5 ; y1 = 4 x2 = 15 ; y2 = -4 m = (-4 - 4)/(15 - (-5)) m = (-8)/(20) m = -2/5

OpenStudy (nurali):

Slope Formula: m=y2-y1/x2-x1